# Exploring Potential Prompt Injection Attacks in Federated Military LLMs and Their Mitigation

### v1

@@ VI Conclusion
In this paper, we present potential prompt injection vulnerabilities within federated military LLMs, pinpointing four key attack scenarios. We have highlighted the wide-ranging impact of these threats in military settings by presenting how secret data can be leaked, free-riders can exploit the system, operations can be disrupted, and false information can spread. In response, we introduce a collaborative human–AI framework which includes perspective of technical and political way, enabling agile defense against rapidly evolving attack patterns. Our findings underscore the necessity of ongoing refinement and coordinated governance, offering a robust foundation for protecting federated military LLMs against evolving prompt injection attacks.

### v2

@@ VI Conclusion
In this work, we present potential prompt injection vulnerabilities within federated military LLMs, pinpointing four key attack scenarios. We highlight the wide-ranging impact of these threats in military settings by presenting how secret data can be leaked, free-riders can exploit the system, operations can be disrupted, and false information can spread. In response, we introduce a collaborative human–AI framework which combines both technical and policy perspectives, enabling agile defense against rapidly evolving attack patterns. Our findings emphasize the need for continual improvement and joint governance to safeguard federated military LLMs against evolving prompt injection attacks.